OpenStudy (anonymous):

The perimeter of a rectangle is 40 feet. The length of the rectangle is 2 feet more than 5 times the width. Find the dimensions of the rectangle.

OpenStudy (anonymous):

Okay so remember the equation for perimeter?

OpenStudy (anonymous):

no, what is it?

OpenStudy (anonymous):

\(\large P_{rectangle}=2l+2w\)

OpenStudy (anonymous):

So we know that \(40 = 2l+2w\) and that \(l = 5w+2\).

OpenStudy (anonymous):

Do you know how to solve for \(l\) and \(w\) here?

OpenStudy (anonymous):

Distribute 2 to 5w and 2 ?

OpenStudy (anonymous):

Well, we substitute \(5w+2\) into the \(l\) of \(40=2l+2w\) to get \(40=2(5w+2) + 2w\)... Then we can distribute the two... yes.

OpenStudy (anonymous):

I got w = 3

OpenStudy (anonymous):

Is that the dimensions of the rectangle?

OpenStudy (anonymous):

Well, now you have to solve for \(l\) by plugging in that \(3\). So we have \(l = 5(3)+2\)

OpenStudy (anonymous):

thanks so much